# Joint Beamforming and Antenna Position Optimization for Fluid Antenna-Assisted MU-MIMO Networks

### v1

@@ IV-D The Decentralized Update of $\underline{\mathbf{W}}$
Proposition 1 (Matrix non-homogeneous transform [30, Corollary 19]).
@@
$\underset{\mathbf{X},\mathbf{Z}\in\mathcal{X}}{\max}\ -\mathtt{tr}\left(\mathbf{X}^{\mathsf{H}}\mathbf{M}\mathbf{X}+2\Re\left(\mathbf{X}^{\mathsf{H}}\left(\mathbf{L}-\mathbf{M}\right)\mathbf{Z}\right)+\mathbf{Z}^{\mathsf{H}}\left(\mathbf{M}-\mathbf{L}\right)\mathbf{Z}\right),$ (55)
in the sense that they achieve identical optimal objective values with identical optimal solutions, where $\mathbf{Z}$ is introduced as an auxiliary variable.
@@
where $f_{\rm NonH}\left(\underline{\mathbf{W}},\underline{\mathbf{\Psi}}\right)$ is given as (58) at the bottom of the page. For simplicity, we define $\underline{\mathbf{\Psi}}=\left\{\mathbf{\Psi}_{k},\forall k\right\}$. Since the problem (57) involves both the auxiliary variables $\underline{\mathbf{\Psi}}$ and the beamforming matrices $\underline{\mathbf{W}}$, we update one set of variables while keeping the other fixed.
@@
Next, we update $\underline{\mathbf{W}}$ while keeping $\underline{\mathbf{\Psi}}$ fixed. By substituting (59) into (58) and setting the first-order derivative of $f_{\rm NonH}\left(\underline{\mathbf{W}},\overline{\underline{\mathbf{W}}}\right)$ w.r.t. $\mathbf{W}_{k}$ to zero, we obtain the closed-form solution to the problem (57), given by
$\mathbf{W}_{k}=\mathbf{Q}_{k}\min\left\{\sqrt{\frac{P_{\max}}{\sum_{j=1}^{K}\mathtt{tr}\left(\mathbf{Q}_{j}^{\mathsf{H}}\mathbf{Q}_{j}\right)}},1\right\}.$ (60)
@@
The matrix $\mathbf{Q}_{k}$ denotes the beamforming matrix without the power constraint, and is given by
$\mathbf{Q}_{k}=$ $\frac{1}{\eta}\left(\sqrt{\alpha_{k}}\overline{\mathbf{H}}_{k}^{\mathsf{H}}\overline{\mathbf{\Phi}}_{k}\left(\mathbf{I}+\overline{\mathbf{\Gamma}}_{k}\right)\right.$
$-\left.\left(\sum_{j=1}^{K}\overline{\mathbf{H}}_{j}^{\mathsf{H}}\overline{\mathbf{\Phi}}_{j}\left(\mathbf{I}+\overline{\mathbf{\Gamma}}_{j}\right)\overline{\mathbf{\Phi}}_{j}^{\mathsf{H}}\overline{\mathbf{H}}_{j}-\eta\mathbf{I}\right)\overline{\mathbf{W}}_{k}\right).$ (61)
Although the per-iteration complexity is significantly reduced by eliminating matrix inversions, more iterations are required for convergence, which may still be time-consuming. Therefore, it is necessary to reduce the number of iterations.
As demonstrated in [31], we can reduce the number of iterations by using momentum methods. Among these methods, Nesterov’s extrapolation strategy [39] extrapolates $\mathbf{W}_{k}$ along the direction defined by the two previous iterations, $\overline{\mathbf{W}}_{k}$ and $\overline{\overline{\mathbf{W}}}_{k}$, to predict its value in the following iteration. This approach has been shown to be effective in this scenario [31]. The extrapolated value $\mathbf{\Upsilon}_{k}$ is defined as
$\mathbf{\Upsilon}_{k}\triangleq\overline{\mathbf{W}}_{k}+\nu_{i}\left(\overline{\mathbf{W}}_{k}-\overline{\overline{\mathbf{W}}}_{k}\right).$ (62)
where $\nu_{i}$ denotes the extrapolation step size in the $i$-th BCA iteration, and is given by
$\nu_{i}=\max\left\{\frac{i-2}{i+1},0\right\}.$ (63)
Using Nesterov’s extrapolation strategy, the matrix $\mathbf{Q}_{k}$ is calculated as
@@
$\mathbf{Q}_{k}=$ $\eta^{-1}\left(\sqrt{\alpha_{k}}\overline{\mathbf{H}}_{k}^{\mathsf{H}}\overline{\mathbf{\Phi}}_{k}\left(\mathbf{I}+\overline{\mathbf{\Gamma}}_{k}\right)\right.$
@@
$-\left.\left(\sum_{j=1}^{K}\overline{\mathbf{H}}_{j}^{\mathsf{H}}\overline{\mathbf{\Phi}}_{j}\left(\mathbf{I}+\overline{\mathbf{\Gamma}}_{j}\right)\overline{\mathbf{\Phi}}_{j}^{\mathsf{H}}\overline{\mathbf{H}}_{j}-\eta\mathbf{I}\right)\mathbf{\Upsilon}_{k}\right).$ (64)
Then, the matrix $\mathbf{W}_{k}$ is obtained by substituting (IV-D) into (60).
$\delta^{\mathtt{Tx}}=\underset{1\leq m\leq M}{\max}\frac{24\pi^{2}}{\lambda^{2}}$ $\sum\nolimits_{k=1}^{K}L_{k}^{t}\left(\left(\sum\nolimits_{t=1}^{K}\Big{\lVert}[\overline{\mathbf{W}}_{t}^{c}]_{m}\Big{\rVert}_{2}\sum\nolimits_{j=1}^{M}\Big{\lVert}[\overline{\mathbf{W}}_{t}]_{j}\Big{\rVert}_{2}+\sum\nolimits_{t=1}^{K}\sum\nolimits_{s=1}^{K}[\overline{\mathbf{W}}_{t}^{c}]_{m}\tilde{\mathbf{W}}_{ts}[\overline{\mathbf{W}}_{s}^{c}]_{m}^{\mathsf{H}}\right)\right.$ (78)
×$\left.\times\lVert\mathbf{\Sigma}_{k}^{\mathsf{H}}\overline{\mathbf{F}}_{k}\overline{\mathbf{\Phi}}_{k}\left(\mathbf{I}+\overline{\mathbf{\Gamma}}_{k}\right)\overline{\mathbf{\Phi}}_{k}^{\mathsf{H}}\overline{\mathbf{F}}_{k}^{\mathsf{H}}\mathbf{\Sigma}_{k}\rVert_{2}+\sqrt{\frac{\alpha_{k}}{L_{k}^{t}}}\Big{\lVert}[\overline{\mathbf{W}}_{k}^{c}]_{m}\left(\mathbf{I}+\overline{\mathbf{\Gamma}}_{k}\right)\overline{\mathbf{\Phi}}_{k}^{\mathsf{H}}\overline{\mathbf{F}}_{k}^{\mathsf{H}}\mathbf{\Sigma}_{k}^{\mathsf{H}}\Big{\rVert}_{2}\right).$
$\delta^{\mathtt{Rx}}_{k}=\underset{1\leq n\leq N}{\max}\frac{24\pi^{2}}{\lambda^{2}}$ $L_{k}^{r}\left(\left(\sum\nolimits_{j=1}^{N}\Big{\lvert}[\overline{\mathbf{\Phi}}_{k}]_{n}\left(\mathbf{I}+\overline{\mathbf{\Gamma}}_{k}\right)[\overline{\mathbf{\Phi}}_{k}]_{j}^{\mathsf{H}}\Big{\rvert}+\sqrt{N}\Big{\lVert}[\overline{\mathbf{\Phi}}_{k}]_{n}\left(\mathbf{I}+\overline{\mathbf{\Gamma}}_{k}\right)\overline{\mathbf{\Phi}}_{k}^{\mathsf{H}}\Big{\rVert}_{2}\right)\right.$ (79)
×$\left.\times\Big{\lVert}\sum\nolimits_{t=1}^{K}\mathbf{\Sigma}_{k}\tilde{\mathbf{G}}_{kt}\tilde{\mathbf{G}}_{kt}^{\mathsf{H}}\mathbf{\Sigma}_{k}^{\mathsf{H}}\Big{\rVert}_{2}+\sqrt{\frac{\alpha_{k}}{L_{k}^{r}}}\Big{\lVert}[\mathbf{\Phi}_{k}]_{n}\left(\mathbf{I}+\overline{\mathbf{\Gamma}}_{k}\right)\tilde{\mathbf{G}}_{kk}^{\mathsf{H}}\mathbf{\Sigma}_{k}^{\mathsf{H}}\Big{\rVert}_{2}\right).$
After obtaining the improved closed-form expression of $\underline{\mathbf{W}}$ given in (60) and (IV-D), the decentralized update of $\underline{\mathbf{W}}$ can be achieved correspondingly. First, we compute $\eta$ distributively. By leveraging the positive semi-definiteness of the matrix $\mathbf{I}+\mathbf{\Gamma}_{k}$, we can decompose it by
$\mathbf{I}+\overline{\mathbf{\Gamma}}_{k}=\mathbf{\Xi}_{k}\sqrt{\mathbf{\Lambda}_{k}}\sqrt{\mathbf{\Lambda}_{k}^{\mathsf{T}}}\mathbf{\Xi}_{k}^{\mathsf{H}},$ (65)
where $\mathbf{\Lambda}_{k}\in\mathbb{R}_{+}^{d\times d}$ is a diagonal matrix of eigenvalues, and $\mathbf{\Xi}_{k}\in\mathbb{C}^{d\times d}$ contains the corresponding eigenvectors. Therefore, $\eta$ can be calculated using the trace property of the Frobenius norm, given by
$\eta=\sqrt{\sum\nolimits_{j=1}^{K}\sum\nolimits_{k=1}^{K}\mathtt{tr}\left(\mathbf{P}_{k}^{\mathsf{H}}\mathbf{P}_{j}\mathbf{P}_{k}^{\mathsf{H}}\mathbf{P}_{j}\right)},$ (66)
where we define $\mathbf{P}_{k}\triangleq\overline{\mathbf{H}}_{k}^{\mathsf{H}}\overline{\mathbf{\Phi}}_{k}\mathbf{\Xi}_{k}\sqrt{\mathbf{\Lambda}_{k}}$, and it is distributively computed at the DUs. Specifically, CU broadcasts $\mathbf{\Lambda}_{k}$ and $\mathbf{\Xi}_{k}$ to all DUs, and each DU first computes $\mathbf{P}_{k}^{c}$ as
$\mathbf{P}_{k}^{c}\triangleq\left(\overline{\mathbf{H}}_{k}^{c}\right)^{\mathsf{H}}\overline{\mathbf{\Phi}}_{k}\mathbf{\Xi}_{k}\sqrt{\mathbf{\Lambda}_{k}}.$ (67)
Then, each DU computes $\left(\mathbf{P}_{k}^{c}\right)^{\mathsf{H}}\mathbf{P}_{j}^{c}$ locally and transmits the result to the CU. Finally, the CU aggregates the results as $\tilde{\mathbf{P}}_{kj}=\sum_{c=1}^{C}\left(\mathbf{P}_{k}^{c}\right)^{\mathsf{H}}\mathbf{P}_{j}^{c}$, and computes $\eta$ as
@@
Note that (68) is equivalent to (66) since $\tilde{\mathbf{P}}_{kj}=\mathbf{P}_{k}^{\mathsf{H}}\mathbf{P}_{j}$.
Next, we use the previously computed $\eta$ to calculate $\mathbf{W}_{k}^{c}$, which is stored distributively at each DU. First, we compute the extrapolated beamforming matrices $\mathbf{\Upsilon}_{k}$, which is stored distributively. At the $c$-th DU, the matrix $\mathbf{\Upsilon}_{k}^{c}$ is computed as
$\mathbf{\Upsilon}_{k}^{c}=\overline{\mathbf{W}}_{k}^{c}+\nu_{i}\left(\overline{\mathbf{W}}_{k}^{c}-\overline{\overline{\mathbf{W}}}_{k}^{c}\right).$ (69)
Then, we compute $\mathbf{Q}_{k}$, which is also stored distributively as $\mathbf{Q}_{k}^{c}$. After each DU locally computes $\overline{\mathbf{H}}_{j}^{c}\mathbf{\Upsilon}_{k}^{c}$ and transmits the result to the CU, the CU aggregates the received $\overline{\mathbf{H}}_{j}^{c}\mathbf{\Upsilon}_{k}^{c}$ to obtain $\tilde{\mathbf{\Upsilon}}_{jk}$ as
$\tilde{\mathbf{\Upsilon}}_{jk}=\sum\nolimits_{c=1}^{C}\overline{\mathbf{H}}_{j}^{c}\mathbf{\Upsilon}_{k}^{c},$ (70)
and broadcasts the result $\tilde{\mathbf{\Upsilon}}_{jk}$ to each DU. Therefore, the calculation of $\mathbf{Q}_{k}^{c}$ is given by
$\mathbf{Q}_{k}^{c}=$ $\eta^{-1}\sqrt{\alpha_{k}}\left(\overline{\mathbf{H}}_{k}^{c}\right)^{\mathsf{H}}\overline{\mathbf{\Phi}}_{k}\left(\mathbf{I}+\overline{\mathbf{\Gamma}}_{k}\right)$
$-\eta^{-1}\sum\nolimits_{j=1}^{K}\left(\overline{\mathbf{H}}_{j}^{c}\right)^{\mathsf{H}}\overline{\mathbf{\Phi}}_{j}\left(\mathbf{I}+\overline{\mathbf{\Gamma}}_{j}\right)\overline{\mathbf{\Phi}}_{j}^{\mathsf{H}}\tilde{\mathbf{\Upsilon}}_{jk}+\mathbf{\Upsilon}_{k}^{c}.$ (71)
Finally, we compute $\mathbf{W}^{c}_{k}$ based on $\mathbf{Q}_{k}^{c}$. We define $\mathtt{tr}(\tilde{\mathbf{Q}}_{k})\triangleq\mathtt{tr}\left(\mathbf{Q}_{k}^{\mathsf{H}}\mathbf{Q}_{k}\right)$ and compute them distributively. Specifically, each DU locally computes $\mathtt{tr}(\left(\mathbf{Q}_{k}^{c}\right)^{\mathsf{H}}\mathbf{Q}_{k}^{c})$ and transmits the result to the CU. The CU then aggregates the received values to compute $\mathtt{tr}(\tilde{\mathbf{Q}}_{k})$ as
$\mathtt{tr}\left(\tilde{\mathbf{Q}}_{k}\right)=\sum\nolimits_{c=1}^{C}\mathtt{tr}\left(\left(\mathbf{Q}_{k}^{c}\right)^{\mathsf{H}}\mathbf{Q}_{k}^{c}\right).$ (72)
Once $\mathtt{tr}(\tilde{\mathbf{Q}}_{k})$ is obtained, it will be broadcasted to all DUs. Finally, $\mathbf{W}^{c}_{k}$ can be computed at the $c$-th DU as
$\mathbf{W}_{k}^{c}=\mathbf{Q}_{k}^{c}\min\left\{\sqrt{\frac{P_{\max}}{\sum_{j=1}^{K}\mathtt{tr}(\tilde{\mathbf{Q}}_{j})}},1\right\}.$ (73)
IV-E Decentralized Update of $\mathbf{T}$ and $\underline{\mathbf{R}}$
To update $\mathbf{T}$ and $\underline{\mathbf{R}}$ distributively, we must compute $\mathbf{D}_{k}^{\mathtt{Tx}}$, $\mathbf{D}_{k}^{\mathtt{Rx}}$, $\delta^{\mathtt{Tx}}$, and $\delta^{\mathtt{Rx}}_{k}$ distributively in advance. The matrix $\mathbf{D}_{k}^{\mathtt{Tx}}$ is stored distributively as $[\mathbf{D}_{k}^{\mathtt{Tx}}]^{c}$, whereas $\mathbf{D}_{k}^{\mathtt{Rx}}$ is stored at the CU. To calculate $\mathbf{D}_{k}^{\mathtt{Tx}}$, the CU broadcasts $\overline{\mathbf{\Gamma}}_{k}$, $\overline{\mathbf{\Phi}}_{k}$, $\overline{\mathbf{F}}_{k}$, and $\tilde{\mathbf{G}}_{kj}$ to each DU. Then, the $c$-th DU computes $[\mathbf{D}_{k}^{\mathtt{Tx}}]^{c}$ as

### v2

@@ V-D Decentralized Update of $\underline{\mathbf{W}}$
Proposition 1 (Matrix non-homogeneous transform [34, Corollary 19]).
@@
$\underset{\mathbf{X},\mathbf{\Psi}\in\mathcal{X}}{\max}\ -\mathtt{tr}\left(\mathbf{X}^{\mathsf{H}}\mathbf{M}\mathbf{X}+2\Re\left(\mathbf{X}^{\mathsf{H}}\left(\mathbf{L}-\mathbf{M}\right)\mathbf{\Psi}\right)+\mathbf{\Psi}^{\mathsf{H}}\left(\mathbf{M}-\mathbf{L}\right)\mathbf{\Psi}\right),$ (58)
in the sense that they achieve identical optimal objective values with identical optimal solutions, where $\mathbf{\Psi}$ is introduced as an auxiliary variable.
@@
where $f_{\rm NonH}\left(\underline{\mathbf{W}},\underline{\mathbf{\Psi}}\right)$ is given as (61) at the top of the next page. For simplicity, we define $\underline{\mathbf{\Psi}}=\left\{\mathbf{\Psi}_{k},\forall k\right\}$. Since the problem (60) involves both the auxiliary variables $\underline{\mathbf{\Psi}}$ and the beamforming matrices $\underline{\mathbf{W}}$, we update one set of variables while keeping the other fixed.
@@
Next, we update $\underline{\mathbf{W}}$ while keeping $\underline{\mathbf{\Psi}}$ fixed. By substituting (62) into (61) and setting the first-order derivative of $f_{\rm NonH}\left(\underline{\mathbf{W}},\overline{\underline{\mathbf{W}}}\right)$ w.r.t. $\mathbf{W}_{k}$ to zero, we obtain the closed-form solution to the problem (60):
$\mathbf{W}_{k}=\mathbf{Q}_{k}\min\Big\{\sqrt{P_{\max}/P_{Q}},1\Big\},$ (63)
@@
where $P_{Q}$ denotes the transmission power with beamforming matrix $\mathbf{Q}_{k}$:
$P_{Q}=\sum\nolimits_{j=1}^{K}\mathtt{tr}\left(\mathbf{Q}_{j}^{\mathsf{H}}\mathbf{Q}_{j}\right),$ (64)
and the matrix $\mathbf{Q}_{k}$ denotes the beamforming matrix without the power constraint:
@@
$\mathbf{Q}_{k}=$ $\eta^{-1}\Big[\sqrt{\alpha_{k}}\overline{\mathbf{H}}_{k}^{\mathsf{H}}\overline{\mathbf{\Phi}}_{k}\left(\mathbf{I}+\overline{\mathbf{\Gamma}}_{k}\right)$
$-\Big(\sum_{j=1}^{K}\overline{\mathbf{H}}_{j}^{\mathsf{H}}\overline{\mathbf{\Phi}}_{j}\left(\mathbf{I}+\overline{\mathbf{\Gamma}}_{j}\right)\overline{\mathbf{\Phi}}_{j}^{\mathsf{H}}\overline{\mathbf{H}}_{j}-\eta\mathbf{I}\Big)\overline{\mathbf{W}}_{k}\Big].$ (65)
Although the per-iteration complexity is significantly reduced by eliminating matrix inversions, more iterations are required for convergence, which may still be time-consuming. Therefore, it is necessary to reduce the number of iterations. As demonstrated in [35], we can reduce the number of iterations by using momentum methods. Among these methods, Nesterov’s extrapolation strategy [50] extrapolates $\mathbf{W}_{k}$ along the direction defined by the two previous iterations, $\overline{\mathbf{W}}_{k}$ and $\overline{\overline{\mathbf{W}}}_{k}$, to predict its value in the following iteration. This approach is effective in this scenario [35]. The extrapolated value $\mathbf{\Upsilon}_{k}$ is defined as
$\mathbf{\Upsilon}_{k}\triangleq\overline{\mathbf{W}}_{k}+\nu_{i}\left(\overline{\mathbf{W}}_{k}-\overline{\overline{\mathbf{W}}}_{k}\right),$ (66)
where $\nu_{i}=\max\left\{(i-2)/(i+1),0\right\}$ represents the extrapolation step size in the $i$-th BCA iteration. Using Nesterov’s extrapolation strategy, the matrix $\mathbf{Q}_{k}$ is calculated as
@@
$\mathbf{Q}_{k}=$ $\eta^{-1}\Big[\sqrt{\alpha_{k}}\overline{\mathbf{H}}_{k}^{\mathsf{H}}\overline{\mathbf{\Phi}}_{k}\left(\mathbf{I}+\overline{\mathbf{\Gamma}}_{k}\right)$
$-\Big(\sum_{j=1}^{K}\overline{\mathbf{H}}_{j}^{\mathsf{H}}\overline{\mathbf{\Phi}}_{j}\left(\mathbf{I}+\overline{\mathbf{\Gamma}}_{j}\right)\overline{\mathbf{\Phi}}_{j}^{\mathsf{H}}\overline{\mathbf{H}}_{j}-\eta\mathbf{I}\Big)\mathbf{\Upsilon}_{k}\Big].$ (67)
Then, the matrix $\mathbf{W}_{k}$ is obtained by substituting (V-D) into (63).
After obtaining the improved closed-form expression of $\underline{\mathbf{W}}$ given in (63) and (V-D), the decentralized update of $\underline{\mathbf{W}}$ can be achieved correspondingly. First, we compute $\eta$ in a distributed manner. Since $\mathbf{I}+\overline{\mathbf{\Gamma}}_{k}\succ\mathbf{0}$, we perform the eigenvalue decomposition (EVD) to obtain $\mathbf{\Lambda}_{k}\in\mathbb{R}_{+}^{d\times d}$, a diagonal matrix of eigenvalues, and $\mathbf{\Xi}_{k}\in\mathbb{C}^{d\times d}$, whose columns are the corresponding eigenvectors. The CU broadcasts $\mathbf{\Lambda}_{k}$ and $\mathbf{\Xi}_{k}$ to all DUs, and each DU computes $\mathbf{P}_{k}^{c}$ as
$\mathbf{P}_{k}^{c}\triangleq(\overline{\mathbf{H}}_{k}^{c})^{\mathsf{H}}\overline{\mathbf{\Phi}}_{k}\mathbf{\Xi}_{k}\sqrt{\mathbf{\Lambda}_{k}}.$ (70)
Then, the value $\tilde{\mathbf{P}}_{kj}\triangleq\mathsf{Mul}(\mathbf{P}_{k}^{c},\mathbf{P}_{j}^{c})$ is calculated similar to the process in Fig. 3, and the CU computes $\eta$ as
@@
Note that (71) is equivalent to (59) and is derived using the trace property of the Frobenius norm.
Next, we use the previously computed $\eta$ to calculate $\mathbf{W}_{k}^{c}$. To avoid repetition of formulas, we refer directly to the centralized equations introduced earlier. The only modification in the decentralized setting is to append a superscript $c$ to the terms $\overline{\mathbf{H}}_{k}$, $\overline{\mathbf{W}}_{k}$, $\overline{\overline{\mathbf{W}}}_{k}$, $\mathbf{\Upsilon}_{k}$, and $\mathbf{Q}_{k}$. First, we compute the extrapolated beamforming matrices $\mathbf{\Upsilon}_{k}^{c}$ at the $c$-th DU by (66). Then, we compute $\tilde{\mathbf{\Upsilon}}_{jk}$ similar to Fig. 3 and compute $\mathbf{Q}_{k}^{c}$ via (V-D). To calculate the value of $P_{Q}$, each DU locally computes $\mathtt{tr}(\left(\mathbf{Q}_{k}^{c}\right)^{\mathsf{H}}\mathbf{Q}_{k}^{c})$ and transmits the result to the CU. The CU then aggregates the received values by $P_{Q}=\sum\nolimits_{k=1}^{K}\sum\nolimits_{c=1}^{C}\mathtt{tr}\left(\left(\mathbf{Q}_{k}^{c}\right)^{\mathsf{H}}\mathbf{Q}_{k}^{c}\right)$. Once $P_{Q}$ is obtained, it is broadcast to all DUs, and $\mathbf{W}^{c}_{k}$ can be computed at the $c$-th DU by (63). The aforementioned process is summarized in Fig. 4, where only the $c$-th DU is shown for brevity.
Figure 4: The decentralized calculation of $\underline{\mathbf{W}}$ under the DBP architecture.
V-E Decentralized Update of $\mathbf{T}$ and $\underline{\mathbf{R}}$